1. A resistor with resistance R = 30 MΩ is connected in series with a capacitor with capacitance 3 μF. What is the


time constant?



2. Determine the equivalent (total) resistance for each of the following circuits below.

1
Expert's answer
2022-03-11T08:33:54-0500

Answer

1.the time constant is given by

"\tau=RC\\=30*10^6*3*10^{-6}\\=900sec"

2.the equivalent (total) resistance for each circuit using

If in series then

"R_T=R+R'+R"+....."

If in parallel

"\frac{1}{R_T}=\frac{1}{R}+\frac{1}{R'}+......."